OpenStudy (gnrfan):

Given the function f(x) = x3 + 2x2 − 3x − 5, what is the resulting function when f(x) is shifted to the right 1 unit? wait a sec for answers below

OpenStudy (gnrfan):

f(x + 1) = x3 + 5x2 + 4x − 5 f(x) + 1 = x3 + 2x2 − 3x − 4 f(x) − 1 = x3 + 2x2 − 3x − 6 f(x − 1) = x3 − x2 − 4x − 1

OpenStudy (gnrfan):

@ganeshie8 @hartnn

OpenStudy (gnrfan):

@Compassionate

OpenStudy (gnrfan):

@paki

OpenStudy (gnrfan):

@perl can you help please

OpenStudy (gnrfan):

@iGreen.

OpenStudy (gnrfan):

do you know igreen?

OpenStudy (anonymous):

I believe subtracting 1 from 'x' in f(x) will shift it to the right..so what should it be?

OpenStudy (gnrfan):

@Legends

OpenStudy (gnrfan):

c or d right?

OpenStudy (anonymous):

Yes, but it will be inside the parenthesis.

OpenStudy (anonymous):

So which one? @gnrfan

OpenStudy (gnrfan):

i think d

OpenStudy (anonymous):

Yep, you got it!

OpenStudy (gnrfan):

yea its D

OpenStudy (anonymous):

\(\bbox [10pt, lime, border:5pt solid black]{\Huge\sf\color{blue}{Welcome\ to\ \color{#00A1FF}{Open}\color{#329932}{Study!!!}}}\) You can give medals by clicking 'Best Response'.

OpenStudy (gnrfan):

can you help with another please

OpenStudy (anonymous):

Sure, just close this and open a new one.

OpenStudy (gnrfan):

@iGreen. can you help again pls

OpenStudy (gnrfan):

k